# A Comprehensive Survey on World Models for Embodied AI

### v1

@@ 4.1 Data Resources
MuJoCo[218] 2012 Continuous control Proprio. Sim - -
CARLA[219] 2017 Driving simulation RGB-D/Seg/LiDAR/Radar/GPS/IMU Sim -
Habitat[220] 2019 Embodied navigation RGB-D/Seg/GPS/Compass Sim -
Isaac Gym[221] 2021 continuous control Proprio. Sim - -
Isaac Lab[222] 2023 Robot learning suites RGB-D/Seg/LiDAR/Proprio. Sim - -
@@
Atari[223] 2013 Discrete-action game RGB/State Sim 55+ Games
DMC[224] 2018 Continuous control RGB/Proprio. Sim 30+ Tasks
Meta-World[225] 2019 Multi-task manipulation RGB/Proprio. Sim 50 tasks
RLBench[226] 2020 Robotic manipulation RGB-D/Seg/Proprio. Sim 100 tasks
nuPlan[227] 2021 Driving planning RGB/LiDAR/Map/Proprio. Real 1.5k hours
LIBERO[228] 2023 Lifelong manipulation RGB/Text/Proprio. Sim 130 tasks
@@
SSv2[229] 2018 Video-action understanding RGB/Text Real 220k videos 169k/24k/27k
nuScenes[230] 2020 Driving perception RGB/LiDAR/Radar/GPS/IMU Real 1k scenes 700/150/150
Waymo[231] 2020 Driving perception RGB/LiDAR Real 1.15k scenes 798/202/150
HM3D[232] 2021 Indoor navigation RGB-D Real 1k scenes 800/100/100
RT-1[233] 2022 Real-robot manipulation RGB/Text Real 130k+ trajectories -
Occ3D[234] 2023 3D occupancy RGB/LiDAR Real 1.9k scenes 600/150/150; 798/202/-
OXE[235] 2024 Cross-embodiment pretraining RGB-D/LiDAR/Text Real 1M+ trajectories -
OpenDV[90] 2024 Driving video pretraining RGB/Text Real 2k+ hours -
VideoMix22M[14] 2025 Video pretraining RGB Real 22M+ samples -
@@
Franka Emika[236] 2022 Manipulation Proprio. Real - -
Unitree Go1 [237] 2021 Quadruped locomotion RGB-D/LiDAR/Proprio. Real - -
Unitree G1[238] 2024 Humanoid manipulation RGB-D/LiDAR/Proprio./Audio Real - -
@@ 4.1.1 Simulation Platforms
  • MuJoCo [218] is a customizable physics engine widely adopted for its efficient robotic simulation of articulated systems and contact dynamics in robotics and control research.
  • NVIDIA Isaac is an E2E, GPU-accelerated simulation stack that comprises Isaac Sim, Isaac Gym [221], and Isaac Lab [222]. It offers photorealistic rendering and large-scale RL capabilities.
  • CARLA [219] is an open-source simulator based on Unreal Engine for urban autonomous driving, providing realistic rendering, diverse sensors, and closed-loop evaluation protocols.
  • Habitat [220] is a high-performance simulator for embodied AI, specializing in photorealistic 3D indoor navigation.
4.1.2 Interactive Benchmarks
@@
  • DeepMind Control (DMC) [224] is a standard MuJoCo-based suite for control tasks, offering a consistent basis for comparing agents that learn from state or pixel-based observations.
  • Atari [223] is a suite of pixel-based, discrete-action games for evaluating agent performance. The Atari100k [239] specifically assesses sample efficiency by limiting interaction to 100k steps.
  • Meta-World [225] is a benchmark for multi-task and meta-RL, featuring $50$ diverse robotic manipulation tasks with a Sawyer arm in MuJoCo under standardized evaluation protocols.
  • RLBench [226] offers $100$ simulated tabletop manipulation tasks with sparse rewards and rich, multi-modal observations, designed to test complex skills and rapid adaptation.
  • LIBERO [228] is a benchmark for lifelong robotic manipulation, providing $130$ procedurally generated tasks and human demonstrations to evaluate sample-efficient and continual learning.
  • nuPlan [227] is a planning benchmark for autonomous driving, using a lightweight closed-loop simulator and over $1\,500\text{\,}\mathrm{h}\text{/}$ of real-world driving logs to evaluate long-horizon performance.
4.1.3 Offline Datasets
@@
  • RT-1 [233] is a real-world dataset for robot learning, collected over $17$ months with $13$ Everyday Robots mobile manipulators. It contains 130 000130\,000 demonstrations spanning more than $700$ tasks, pairing language instructions and image observations with discretized 11-DoF actions for the arm and mobile base.
  • Open X-Embodiment (OXE) [235] is a corpus aggregating $60$ sources from $21$ institutions, spanning $22$ robot embodiments, $527$ skills, and over one million trajectories in a unified format for cross-embodiment training. Models trained on OXE demonstrate strong transfer beyond single-robot baselines, underscoring the effectiveness of cross-platform data sharing.
@@
  • nuScenes [230] is a large-scale multimodal driving dataset with a 360-degree sensor suite comprising six cameras, five radars, one LiDAR, and GPS/IMU. It contains 1 0001\,000 twenty-second scenes collected in Boston and Singapore with dense 3D annotations for $23$ classes and HDMaps, providing a core benchmark for multimodal fusion and long-horizon prediction.
  • Waymo [231] is a multimodal autonomous driving benchmark with 1 1501\,150 twenty-second scenes at $10\text{\,}\mathrm{Hz}\text{/}$ from San Francisco, Phoenix, and Mountain View. It includes five LiDARs and five cameras, with about $12$ million 3D and 2D annotations, making it a large-scale resource for modeling traffic dynamics.
  • Occ3D [234] defines 3D occupancy prediction from surround-view images, providing voxel labels that distinguish free, occupied, and unobserved states. Occ3D-nuScenes contains about 40 00040\,000 frames at $0.4\text{\,}\mathrm{m}\text{/}$ resolution, while Occ3D-Waymo offers about 200 000200\,000 frames at $0.05\text{\,}\mathrm{m}\text{/}$. This voxel-level supervision enables holistic scene understanding beyond bounding boxes.
  • Something-Something v2 (SSv2) [229] is a video dataset for fine-grained action understanding. It contains 220 847220\,847 clips across $174$ categories, collected from crowd workers following textual prompts (e.g., Putting something into something) with splits of 168 913168\,913 train, 24 77724\,777 val, and 27 15727\,157 test videos.
  • OpenDV [90] is the largest large-scale video-text dataset for autonomous driving, proposed by GenAD, which supports video prediction and world-model pretraining. It contains 2 0592\,059 hours and $65.1$ million frames from YouTube and seven public datasets, covering over $40$ countries and $244$ cities. The dataset provides command and context annotations to enable language- and action-conditioned prediction and planning.
  • VideoMix22M [14] is a large-scale dataset introduced with V-JEPA 2 for self-supervised pretraining. It scales from $2$ million to $22$ million samples, drawn from YT-Temporal-1B [240], HowTo100M [241], Kinetics [242], SSv2, and ImageNet [243]. The largest source, YT-Temporal-1B, is curated with retrieval-based filtering to suppress noise, while ImageNet images are converted into static video clips for consistency.
4.1.4 Real-world Robot Platforms
@@
  • Franka Emika [236] is a 7-DoF collaborative robot arm with joint torque sensors for precise force control. Through the control interface, it supports $1\text{\,}\mathrm{kHz}\text{/}$ torque control for contact-rich tasks, while its ROS integration makes it a versatile platform.
  • Unitree Go1 [237] is a cost-effective and widely adopted quadrupedal robot equipped with a panoramic depth-sensing suite, onboard computing of 1.5 TFLOPS, and a maximum speed of $4.7\text{\,}\mathrm{m}\text{/}\mathrm{s}$, establishing it as a standard platform for locomotion and embodied-AI research.
@@
4.2 Metrics
Metrics evaluate the capability of world models to capture dynamics, generalize to unseen scenarios, and scale with additional resources. We organize them into three abstraction levels: §4.2.1 Pixel Prediction Quality, §4.2.2 State-level Understanding, and §4.2.3 Task Performance, representing a progression from low-level signal fidelity to high-level goal attainment.
4.2.1 Pixel Generation Quality
@@
Fréchet Inception Distance (FID) [244].  FID is a metric for assessing the realism and diversity of generated images. It compares real and generated image distributions in the feature space of an ImageNet-pretrained Inception-v3 [245], modeling embeddings as Gaussians with means $\bm{\mu}_{x},\bm{\mu}_{y}$ and covariances $\bm{\Sigma}_{x},\bm{\Sigma}_{y}$. Defined as
@@
Fréchet Video Distance (FVD) [246].  FVD extends FID to videos, evaluating both per-frame quality and temporal consistency. It replaces the image-based Inception network with an I3D [247] pretrained on Kinetics-400 [248]. Using the same Fréchet distance as Eq. (6) on motion-aware features, FVD yields a holistic video quality score. A lower value indicates a closer alignment of distributions in appearance and dynamics while penalizing temporal artifacts like unnatural motion or flickering.
Structural Similarity Index Measure (SSIM) [249].  SSIM is a perceptual metric for image quality that compares luminance, contrast, and structure between a generated image and its reference. For two patches $x$ and $y$ with means $\bm{\mu}_{x}$, $\bm{\mu}_{y}$, variances $\bm{\Sigma}_{x}^{2},\bm{\Sigma}_{y}^{2}$, and covariance $\bm{\Sigma}_{xy}$, SSIM is defined as
@@
Peak Signal-to-Noise Ratio (PSNR) [250].  PSNR measures pixel-wise distortion between a reconstruction and its reference. Let the mean-squared error (MSE) over $N$ pixels be
@@
Learned Perceptual Image Patch Similarity (LPIPS) [251].  LPIPS is a metric that correlates with human judgments by comparing features extracted from pretrained networks. Let $\hat{f}^{l}_{x}$ and $\hat{f}^{l}_{y}$ denote the unit-normalized activations at layer $l$ for inputs $x$ and $y$, and $w_{l}$ the channel-wise weights. LPIPS is defined as
@@
VBench [252].  VBench is a comprehensive benchmark for video generation that assesses performance across 16 dimensions grouped into two categories: Video Quality (e.g., subject consistency, motion smoothness) and Video-Condition Consistency (e.g., object class, human action). It provides carefully curated prompt suites and large-scale human preference annotations to ensure strong perceptual alignment, thereby enabling fine-grained evaluation of model capabilities and limitations.
4.2.2 State-level Understanding
@@
Chamfer Distance (CD) [253].  CD quantifies geometric similarity between a prediction $S_{1}$ and ground truth $S_{2}$ by summing squared nearest-neighbor distances across the two sets:
@@ 5.1 Pixel Generation
MagicDrive3D[84] arXiv’24 $224\times 400$ 20.7 164.7
Delphi[86] arXiv’24 $512\times 512$ 15.1 113.5
Drive-WM[88] CVPR’24 $192\times 384$ 15.8 122.7
GenAD[90] CVPR’24 $256\times 448$ 15.4 184.0
DriveDreamer[91] ECCV’24 $128\times 192$ 52.6 452.0
Vista[96] NeurIPS’24 $576\times 1024$ 6.9 89.4
DrivePhysica[214] arXiv’24 $256\times 448$ 4.0 38.1
DrivingWorld[133] arXiv’24 $512\times 1024$ 7.4 90.9
DriveDreamer-2[97] AAAI’25 $256\times 448$ 11.2 55.7
UniFuture[206] arXiv’25 $320\times 576$ 11.8 99.9
MiLA[189] arXiv’25 $360\times 640$ 4.1 14.9
GeoDrive[170] arXiv’25 $480\times 720$ 4.1 61.6
LongDWM[188] arXiv’25 $480\times 720$ 12.3 102.9
MaskGWM[104] CVPR’25 $288\times 512$ 8.9 65.4
GEM[102] CVPR’25 $576\times 1024$ 10.5 158.5
Epona[148] ICCV’25 $512\times 1024$ 7.5 82.8
STAGE[198] IROS’25 $512\times 768$ 11.0 242.8
DriVerse[109] ACMMM’25 $480\times 832$ 18.2 95.2
5.2 Scene Understanding

### v2

@@ IV-A Data Resources
MuJoCo[todorov2012mujoco] 2012 Continuous control Proprio. Sim - -
CARLA[dosovitskiy2017carla] 2017 Driving simulation RGB-D/Seg/LiDAR/Radar/GPS/IMU Sim -
Habitat[savva2019habitat] 2019 Embodied navigation RGB-D/Seg/GPS/Compass Sim -
Isaac Gym[makoviychuk2021isaac] 2021 continuous control Proprio. Sim - -
Isaac Lab[mittal2023orbit] 2023 Robot learning suites RGB-D/Seg/LiDAR/Proprio. Sim - -
@@
Atari[bellemare2013arcade] 2013 Discrete-action game RGB/State Sim 55+ Games
DMC[tassa2018deepmind] 2018 Continuous control RGB/Proprio. Sim 30+ Tasks
Meta-World[yu2020meta] 2019 Multi-task manipulation RGB/Proprio. Sim 50 tasks
RLBench[james2020rlbench] 2020 Robotic manipulation RGB-D/Seg/Proprio. Sim 100 tasks
nuPlan[caesar2021nuplan] 2021 Driving planning RGB/LiDAR/Map/Proprio. Real 1.5k hours
LIBERO[liu2023libero] 2023 Lifelong manipulation RGB/Text/Proprio. Sim 130 tasks
@@
SSv2[goyal2017something] 2018 Video-action understanding RGB/Text Real 220k videos 169k/24k/27k
nuScenes[caesar2020nuscenes] 2020 Driving perception RGB/LiDAR/Radar/GPS/IMU Real 1k scenes 700/150/150
Waymo[sun2020scalability] 2020 Driving perception RGB/LiDAR Real 1.15k scenes 798/202/150
HM3D[ramakrishnan2021habitatmatterport] 2021 Indoor navigation RGB-D Real 1k scenes 800/100/100
RT-1[brohan2022rt] 2022 Real-robot manipulation RGB/Text Real 130k+ trajectories -
Occ3D[tian2023occ3d] 2023 3D occupancy RGB/LiDAR Real 1.9k scenes 600/150/150; 798/202/-
OXE[o2024open] 2024 Cross-embodiment pretraining RGB-D/LiDAR/Text Real 1M+ trajectories -
OpenDV[yang2024generalized] 2024 Driving video pretraining RGB/Text Real 2k+ hours -
VideoMix22M[assran2025v] 2025 Video pretraining RGB Real 22M+ samples -
@@
Franka Emika[haddadin2022franka] 2022 Manipulation Proprio. Real - -
Unitree Go1 [unitree_go1_web] 2021 Quadruped locomotion RGB-D/LiDAR/Proprio. Real - -
Unitree G1[unitree_g1_web] 2024 Humanoid manipulation RGB-D/LiDAR/Proprio./Audio Real - -
@@ IV-A1 Simulation Platforms
  • MuJoCo [todorov2012mujoco] is a customizable physics engine widely adopted for its efficient robotic simulation of articulated systems and contact dynamics in robotics and control research.
  • NVIDIA Isaac is an E2E, GPU-accelerated simulation stack that comprises Isaac Sim, Isaac Gym [makoviychuk2021isaac], and Isaac Lab [mittal2023orbit]. It offers photorealistic rendering and large-scale RL capabilities.
  • CARLA [dosovitskiy2017carla] is an open-source simulator based on Unreal Engine for urban autonomous driving, providing realistic rendering, diverse sensors, and closed-loop evaluation protocols.
  • Habitat [savva2019habitat] is a high-performance simulator for embodied AI, specializing in photorealistic 3D indoor navigation.
IV-A2 Interactive Benchmarks
@@
  • DeepMind Control (DMC) [tassa2018deepmind] is a standard MuJoCo-based suite for control tasks, offering a consistent basis for comparing agents that learn from state or pixel-based observations.
  • Atari [bellemare2013arcade] is a suite of pixel-based, discrete-action games for evaluating agent performance. The Atari100k [kaiser2020model] specifically assesses sample efficiency by limiting interaction to 100k steps.
  • Meta-World [yu2020meta] is a benchmark for multi-task and meta-RL, featuring $50$ diverse robotic manipulation tasks with a Sawyer arm in MuJoCo under standardized evaluation protocols.
  • RLBench [james2020rlbench] offers $100$ simulated tabletop manipulation tasks with sparse rewards and rich, multi-modal observations, designed to test complex skills and rapid adaptation.
  • LIBERO [liu2023libero] is a benchmark for lifelong robotic manipulation, providing $130$ procedurally generated tasks and human demonstrations to evaluate sample-efficient and continual learning.
  • nuPlan [caesar2021nuplan] is a planning benchmark for autonomous driving, using a lightweight closed-loop simulator and over $1\,500\text{\,}\mathrm{h}\text{/}$ of real-world driving logs to evaluate long-horizon performance.
IV-A3 Offline Datasets
@@
  • RT-1 [brohan2022rt] is a real-world dataset for robot learning, collected over $17$ months with $13$ Everyday Robots mobile manipulators. It contains 130 000130\,000 demonstrations spanning more than $700$ tasks, pairing language instructions and image observations with discretized 11-DoF actions for the arm and mobile base.
  • Open X-Embodiment (OXE) [o2024open] is a corpus aggregating $60$ sources from $21$ institutions, spanning $22$ robot embodiments, $527$ skills, and over one million trajectories in a unified format for cross-embodiment training. Models trained on OXE demonstrate strong transfer beyond single-robot baselines, underscoring the effectiveness of cross-platform data sharing.
@@
  • nuScenes [caesar2020nuscenes] is a large-scale multimodal driving dataset with a 360-degree sensor suite comprising six cameras, five radars, one LiDAR, and GPS/IMU. It contains 1 0001\,000 twenty-second scenes collected in Boston and Singapore with dense 3D annotations for $23$ classes and HDMaps, providing a core benchmark for multimodal fusion and long-horizon prediction.
  • Waymo [sun2020scalability] is a multimodal autonomous driving benchmark with 1 1501\,150 twenty-second scenes at $10\text{\,}\mathrm{Hz}\text{/}$ from San Francisco, Phoenix, and Mountain View. It includes five LiDARs and five cameras, with about $12$ million 3D and 2D annotations, making it a large-scale resource for modeling traffic dynamics.
  • Occ3D [tian2023occ3d] defines 3D occupancy prediction from surround-view images, providing voxel labels that distinguish free, occupied, and unobserved states. Occ3D-nuScenes contains about 40 00040\,000 frames at $0.4\text{\,}\mathrm{m}\text{/}$ resolution, while Occ3D-Waymo offers about 200 000200\,000 frames at $0.05\text{\,}\mathrm{m}\text{/}$. This voxel-level supervision enables holistic scene understanding beyond bounding boxes.
  • Something-Something v2 (SSv2) [goyal2017something] is a video dataset for fine-grained action understanding. It contains 220 847220\,847 clips across $174$ categories, collected from crowd workers following textual prompts (e.g., Putting something into something) with splits of 168 913168\,913 train, 24 77724\,777 val, and 27 15727\,157 test videos.
  • OpenDV [yang2024generalized] is the largest large-scale video-text dataset for autonomous driving, proposed by GenAD, which supports video prediction and world-model pretraining. It contains 2 0592\,059 hours and $65.1$ million frames from YouTube and seven public datasets, covering over $40$ countries and $244$ cities. The dataset provides command and context annotations to enable language- and action-conditioned prediction and planning.
  • VideoMix22M [assran2025v] is a large-scale dataset introduced with V-JEPA 2 for self-supervised pretraining. It scales from $2$ million to $22$ million samples, drawn from YT-Temporal-1B [zellers2022merlot], HowTo100M [miech2019howto100m], Kinetics [carreira2019short], SSv2, and ImageNet [deng2009imagenet]. The largest source, YT-Temporal-1B, is curated with retrieval-based filtering to suppress noise, while ImageNet images are converted into static video clips for consistency.
IV-A4 Real-world Robot Platforms
@@
  • Franka Emika [haddadin2022franka] is a 7-DoF collaborative robot arm with joint torque sensors for precise force control. Through the control interface, it supports $1\text{\,}\mathrm{kHz}\text{/}$ torque control for contact-rich tasks, while its ROS integration makes it a versatile platform.
  • Unitree Go1 [unitree_go1_web] is a cost-effective and widely adopted quadrupedal robot equipped with a panoramic depth-sensing suite, onboard computing of 1.5 TFLOPS, and a maximum speed of $4.7\text{\,}\mathrm{m}\text{/}\mathrm{s}$, establishing it as a standard platform for locomotion and embodied-AI research.
@@
IV-B Metrics
Metrics evaluate the capability of world models to capture dynamics, generalize to unseen scenarios, and scale with additional resources. We organize them into three abstraction levels: §IV-B1 Pixel Prediction Quality, §IV-B2 State-level Understanding, and §IV-B3 Task Performance, representing a progression from low-level signal fidelity to high-level goal attainment.
IV-B1 Pixel Generation Quality
@@
Fréchet Inception Distance (FID) [heusel2017gans].  FID is a metric for assessing the realism and diversity of generated images. It compares real and generated image distributions in the feature space of an ImageNet-pretrained Inception-v3 [szegedy2016rethinking], modeling embeddings as Gaussians with means $\bm{\mu}_{x},\bm{\mu}_{y}$ and covariances $\bm{\Sigma}_{x},\bm{\Sigma}_{y}$. Defined as
@@
Fréchet Video Distance (FVD) [unterthiner2018towards].  FVD extends FID to videos, evaluating both per-frame quality and temporal consistency. It replaces the image-based Inception network with an I3D [carreira2017quo] pretrained on Kinetics-400 [kay2017kinetics]. Using the same Fréchet distance as Eq. (6) on motion-aware features, FVD yields a holistic video quality score. A lower value indicates a closer alignment of distributions in appearance and dynamics while penalizing temporal artifacts like unnatural motion or flickering.
Structural Similarity Index Measure (SSIM) [wang2004image].  SSIM is a perceptual metric for image quality that compares luminance, contrast, and structure between a generated image and its reference. For two patches $x$ and $y$ with means $\bm{\mu}_{x}$, $\bm{\mu}_{y}$, variances $\bm{\Sigma}_{x}^{2},\bm{\Sigma}_{y}^{2}$, and covariance $\bm{\Sigma}_{xy}$, SSIM is defined as
@@
Peak Signal-to-Noise Ratio (PSNR) [hore2010image].  PSNR measures pixel-wise distortion between a reconstruction and its reference. Let the mean-squared error (MSE) over $N$ pixels be
@@
Learned Perceptual Image Patch Similarity (LPIPS) [zhang2018unreasonable].  LPIPS is a metric that correlates with human judgments by comparing features extracted from pretrained networks. Let $\hat{f}^{l}_{x}$ and $\hat{f}^{l}_{y}$ denote the unit-normalized activations at layer $l$ for inputs $x$ and $y$, and $w_{l}$ the channel-wise weights. LPIPS is defined as
@@
VBench [huang2024vbench].  VBench is a comprehensive benchmark for video generation that assesses performance across 16 dimensions grouped into two categories: Video Quality (e.g., subject consistency, motion smoothness) and Video-Condition Consistency (e.g., object class, human action). It provides carefully curated prompt suites and large-scale human preference annotations to ensure strong perceptual alignment, thereby enabling fine-grained evaluation of model capabilities and limitations.
IV-B2 State-level Understanding
@@
Chamfer Distance (CD) [fan2017point].  CD quantifies geometric similarity between a prediction $S_{1}$ and ground truth $S_{2}$ by summing squared nearest-neighbor distances across the two sets:
@@ V-A Pixel Generation
MagicDrive3D[gao2024magicdrive3d] arXiv’24 $224\times 400$ 20.7 164.7
Delphi[ma2024unleashing] arXiv’24 $512\times 512$ 15.1 113.5
Drive-WM[wang2024driving] CVPR’24 $192\times 384$ 15.8 122.7
GenAD[yang2024generalized] CVPR’24 $256\times 448$ 15.4 184.0
DriveDreamer[wang2024drivedreamer] ECCV’24 $128\times 192$ 52.6 452.0
Vista[gao2024vista] NeurIPS’24 $576\times 1024$ 6.9 89.4
DrivePhysica[yang2024physical] arXiv’24 $256\times 448$ 4.0 38.1
DrivingWorld[hu2024drivingworld] arXiv’24 $512\times 1024$ 7.4 90.9
DriveDreamer-2[zhao2025drivedreamer] AAAI’25 $256\times 448$ 11.2 55.7
UniFuture[liang2025seeing] arXiv’25 $320\times 576$ 11.8 99.9
MiLA[wang2025mila] arXiv’25 $360\times 640$ 4.1 14.9
GeoDrive[chen2025geodrive] arXiv’25 $480\times 720$ 4.1 61.6
LongDWM[wang2025longdwm] arXiv’25 $480\times 720$ 12.3 102.9
MaskGWM[ni2025maskgwm] CVPR’25 $288\times 512$ 8.9 65.4
GEM[hassan2025gem] CVPR’25 $576\times 1024$ 10.5 158.5
Epona[zhang2025epona] ICCV’25 $512\times 1024$ 7.5 82.8
STAGE[wang2025stage] IROS’25 $512\times 768$ 11.0 242.8
DriVerse[li2025driverse] ACMMM’25 $480\times 832$ 18.2 95.2
V-B Scene Understanding